OpenStudy (anonymous):

Use the distributive property to rewrite the expression and solve. 8*63=8*_______+8*3=________ The *'s are multiply lol

OpenStudy (bohotness):

Use the distributive property to rewrite the expression and solve. 8 * 63 = 8*60 + 8 * 3 = 480 + 24 .

OpenStudy (anonymous):

I dont get it lol